OpenStudy (tomektomeknyc):

(log 4 64) / (2 log 4 √4) is equal to a) log 4 4 b) 4 c) 2 log 4 4 d) 3

OpenStudy (campbell_st):

well start by writing the values in terms of the bases \[64 = 4^3\] \[\sqrt{4} = 4^{\frac{1}{2}}\] so your question is now and applying a log law to the numerator to manage the 2 \[\frac{\log_4(64)}{2\log_{4} \sqrt{4}} = \frac{\log_{4}(4^3)}{\log_{4}(4^{\frac{1}{2}})^2}\] so you are now being asked to simplify \[\frac{\log_{4}(4^3)}{\log_{4}(4^1)}\] all you now need to do is know \[\log_{a}(a^x) = x\] which says, then the base of the log and base of the exponent are the same tit can be simplified to just the power... hope it helps

OpenStudy (tomektomeknyc):

Thank you

OpenStudy (tomektomeknyc):

So I think it is answer c)

OpenStudy (campbell_st):

nope... applied the last law I listed \[\log_{4}(4^3) = 3\] and \[\log_{4}(4^1) = 1\] so you fraction becomes \[\frac{\log_{4}(4^3)}{\log_{4}(4^1)} = \frac{3}{1} = 3\]

OpenStudy (tomektomeknyc):

Thank you once again!